OpenStudy (thefluffmuncher):

A single tree produces about 2.6*10^2 lb of oxygen each year. The Amazon rainforest has about 3.9*10^11 trees. About how many pounds of oyxgen are produced in the rainforest each day? 2.8*10^ lb

OpenStudy (thefluffmuncher):

can someone please explain this to me

OpenStudy (thefluffmuncher):

@Destinymasha

OpenStudy (thefluffmuncher):

@dumbcow

OpenStudy (thefluffmuncher):

@Ashleyisakitty

OpenStudy (thefluffmuncher):

@Preetha

OpenStudy (thefluffmuncher):

@sammixboo

OpenStudy (thefluffmuncher):

@YanaSidlinskiy

OpenStudy (thefluffmuncher):

@pgpilot326

OpenStudy (triciaal):

this a rate problem and need to do conversions

OpenStudy (thefluffmuncher):

i was doing scientific notation

OpenStudy (triciaal):

1 tree = 2.6 * 10^2 lbs oxygen so 3.9* 10^11 times 1 tree= 3.9*10^11 * 2.6 * 10^2 lbs oxygen for the year

OpenStudy (triciaal):

yes it would be 3.9*2.6 * 10^(11+2 ) =

OpenStudy (triciaal):

this is produced in 1 year and the question asks for the day so you have the amount for 365 days and you need to know how much this is per day

OpenStudy (thefluffmuncher):

im need to find the missing exponent